OpenStudy (anonymous):

Solve each system of equations by substitution 2y-6z=-30 -y-z=-1

OpenStudy (nurali):

2y-6z=-30---------(1) -y-z=-1---------(2) in equation (2) -y-z=-1 -y=-1+z -y=z-1 y=-(z-1) y=1-z----------(3) put equation (3) in equation (1) 2(1-z)-6z=-30 2-2z-6z=-30 2-8z=-30 -8z=-32 z=4 put z in equation (3) y=1-4 y=-3 Solution is (-3,4)